-5X -10 = 2 - (X +4): -5X -10 = 2 - (X +4)
Answer by stanbon(19705)

-5X -10 = 2 - (X +4)
-5x - 10 = 2 -x -4
-5x - 10 = -x -2
4x = -8
x = -2
===========
Cheers,
Stan H.